# Measuring Social Norms of Large Language Models

### v2

@@ Appendix D Summary of Skills
Subject: Social Studies Skill: causes-of-the-american-revolution-introduction-to-mercantilism-and-the-navigation-acts Description: To increase Great Britain’s wealth, Parliament passed laws that gave each part of the British Empire a different economic role: Colonists would send raw materials to people in Great Britain. Raw materials, such as raw iron and wool, are directly from nature. People in Great Britain would turn the raw materials into manufactured goods, or finished products, such as fabric and tools. British merchants could sell these manufactured goods back to the colonies or to other countries. Complete the text. Parlaiment wanted the country to earn more money from selling exports than it spent buying imports. The raw materials that Great Britain imported from the colonies were ___ the manufactured goods that Great Britain exported for colonists to purchase. Parliament’s policies were part of what created the system of Triangular Trade. Choices: [(a) more valuable than, (b) as valuable as, (c) less valuable than] Answer index: 2
Subject: Social Studies Skill: the-revolutionary-war-preparing-for-war Description: After 1763, the British Parliament began to tax the Thirteen Colonies more heavily to help pay off debt from the French and Indian War. Parliament also restricted colonial trade. Many colonists resented these new taxes and restrictions. After the passage of the Stamp Act in 1765, which required Americans to pay a tax on all paper they used, an American colonist drew a cartoon with this caption: The moral is, that the colonies may be ruined, but that Britain would thereby be maimed. maimed:permanently wounded What is the author most likely arguing in the caption? Choices: [(a) The colonies are not a part of the British Empire., (b) Great Britain is harming itself by taxing the colonies., (c) Great Britain’s taxation policies are good for the colonies., (d) The colonies will be better off if they declare independence from Great Britain.] Answer index: 1
Subject: Social Studies Skill: u-s-mexican-war Description: By the 1840s, belief in Manifest Destiny had become quite popular in the United States. So, support for expanding the territory of the U.S. was an important issue in the 1844 presidential election. This election was won by James K. Polk, the candidate of the Democratic Party. Read the following passage adapted from the 1844 Democratic Party platform, or list of the party’s goals. Then complete the sentence below. Our title to the whole of the Territory of Oregon is clear and unquestionable, and the annexation of Texas to the United States at the earliest moment possible is important to American progress. annexation:addition Complete the sentence. According to the platform, the Democratic Party ___ Manifest Destiny. Choices: [(a) supported, (b) opposed, (c) neither supported nor opposed] Answer index: 0
Subject: Social Studies Skill: the-constitution-amendments Description: When the Constitution was written, many Americans were afraid that it made the federal government too powerful. So, ten amendments were added to the Constitution to describe and protect the rights of individual people. These first ten amendments are called the Bill of Rights. The following passage is from a letter written by Thomas Jefferson to James Madison in 1787, before the Bill of Rights had been approved. Read the passage. Then answer the question below. A bill of rights is what the people are entitled to . . . and [it is] what no just government should refuse [to write down]. just:fair Based on the passage, what was Jefferson’s attitude toward a bill of rights? Choices: [(a) Governments should protect people’s rights, but those rights don’t need to be written down., (b) Writing down a bill of rights doesn’t do anything, because governments can just ignore it., (c) People deserve to have a bill of rights that describes and protects their individual rights.] Answer index: 2
@@
Subject: Social Studies Skill: the-revolutionary-war-struggle-for-independence Description: As the Continental Army fought the British, the Second Continental Congress sent an Olive Branch Petition to King George III in July 1775. The petition stated that the colonies would stop fighting with Great Britain under certain conditions. Which of the following was one of those conditions? Choices: [(a) if a new colony was established on the western frontier, (b) if Great Britain repealed certain laws and taxes, (c) if the colonies could have their own king and Parliament, (d) if King George III stepped down and a new king was named] Answer index: 1
Subject: Social Studies Skill: the-end-of-reconstruction-and-the-rise-of-jim-crow Description: Some white Southerners believed deeply that they should be shown more respect than African American people. They did not like the Republicans’ plans to protect African American people’s rights. Sometimes their anger at the changes to Southern society turned violent. The passage below comes from an investigation into the murder of an African American man named Johnson Stuart. The man’s brother is describing what happened before the murder. Read the passage. Then answer the question below. Question—Did you ever hear of any threats being made against him before he was killed? . . . State of South Carolina, Evidence taken by the Committee of Investigation of the Third Congressional District, 1870. According to this passage, why did some white Southerners want to kill Johnson Stuart? Choices: [(a) They thought he disrespected white people by treating all the voters the same., (b) They thought he was cheating by paying people to vote for him for sheriff., (c) They thought he had stolen money and supplies from the sheriff’s office in the courthouse., (d) They thought he was violating the Constitution by not allowing African Americans to vote.] Answer index: 0
Subject: Social Studies Skill: pre-contact-native-peoples-introduction Description: For many years, most archaeologists agreed that the first humans who came to North America walked across the land bridge around 12,000 years ago. However, new evidence has been discovered that makes some archaeologists believe people may have traveled in a different way. The text below describes two pieces of evidence that archaeologists use to support a different theory of migration to North America. Look at the text. Then complete the sentence below. Hunter-gatherers lived in the islands of Japan 20,000 years ago. They used resources from the ocean to survive. People on the coast of North America used small boats to travel along the coast more than 10,000 years ago. The text below describes a new theory about how humans arrived in North America. Use the evidence above to complete the text. According to the new theory, the first humans may have traveled from Asia to North America ___. They may have ___ along the coast of the land bridge. Choices: [(a) 100,000 years ago . . . walked, (b) using small boats . . . sailed] Answer index: 1
Subject: Social Studies Skill: the-reformation Description: Read two of Martin Luther’s statements criticizing the Catholic Church. Then answer the question below. It is certain that when money clinks in the [Catholic Church’s] money chest, greed…can be increased. William R. Russell, translator, The Ninety-Five Theses and Other Writings. Copyright 2017 by Penguin Classics. Based on the statements, what was one reason why many people were unhappy with the Catholic Church? Choices: [(a) They thought that the Catholic Church had become too big., (b) They felt their religious community was too small., (c) They believed the Catholic Church should collect more money., (d) They thought the Catholic Church was too rich and greedy.] Answer index: 3
Subject: Social Studies Skill: rights-and-responsibilities-of-active-citizenship Description: American presidents often speak about the rights, duties, and responsibilities of American citizens. Read the quotations from two former presidents. Then answer the question below. Ask not what your country can do for you; ask what you can do for your country. Ronald Reagan, Veterans Day Proclamation, October 26, 1981 Based on the quotations above, which of the following statements would both presidents agree with? Choices: [(a) The country should not expect citizens to do things for their communities., (b) Citizens should expect the government to do everything for them., (c) Active and engaged citizens are an important part of American democracy.] Answer index: 2
@@
Subject: Social Studies Skill: the-age-of-exploration-origins Description: In the 1400s, Europe was made up of many independent states. These states often competed with each other for resources, including imported goods from Africa and Asia. Imagine you are a European ruler in the 1400s. What could you do as a ruler to gain a competitive advantage, or an upper hand, over other European states? Choices: [(a) try to find new trade routes to Africa and Asia, (b) ban any goods that came from Africa or Asia, (c) rely on other European states to sell goods from Africa and Asia to you] Answer index: 0
@@
Subject: Social Studies Skill: industrialization-in-the-gilded-age Description: Industrialization brought dramatic changes to Americans’ daily lives. Read a modern scholar’s description of those changes. Then answer the question below. Within a few decades, urban American homes became networked. . . . Instead of relying on candles and [fuel] carried into the home, each home was connected to the electricity network that provided electric light . . . Instead of relying on . . . outhouses and cesspools, each home was gradually connected to two more networks, one bringing in a supply of clean running water and the other taking waste out into sewers. Houses of the rich after 1880 and of the working class after 1910 were increasingly supplied with central heating. Robert J. Gordon, The Rise and Fall of American Growth, Princeton University Press, 2016. What is the main idea of the passage? Choices: [(a) As the U.S. industrialized, fewer homes had electricity and central heating., (b) Industrialization gave many Americans much more comfortable lives., (c) Electric lighting and sewers made homes dirtier and more dangerous., (d) Factories became very modern, but regular people’s houses did not change.] Answer index: 1
@@
Subject: Social Studies Skill: causes-of-the-american-revolution-politics-and-society Description: Andrew Burnaby, an English minister, traveled around the Thirteen Colonies during the French and Indian War. Based on his observations, Burnaby argued that America would never be independent of Great Britain. Read Burnaby’s argument. Then answer the question below. The difference of character, of manners, of religion, of interest, of the different colonies [is so great that if they were] left to themselves there would soon be a civil war. Why did Burnaby think that America would never be a nation independent from Great Britain? Choices: [(a) The colonies were united in their support of Great Britain., (b) Colonial businesses needed British customers., (c) There were too many differences between the colonies., (d) The colonies signed a pledge to Great Britain after the French and Indian War.] Answer index: 2
Subject: Social Studies Skill: new-england-colonies-founding-and-government Description: Members of a religious group called the Puritans founded the Massachusetts Bay Colony. Unlike the Pilgrims, the Puritans did not want to leave the Church of England. Instead, they wanted to change the Church of England according to their religious beliefs. The first governor of the Massachusetts Bay Colony was a Puritan named John Winthrop. The passage below comes from a speech that Winthrop gave to the Puritans on their voyage to New England. Read the passage. Then answer the question below. For we must consider that we shall be as a city upon a hill. The eyes of all people are upon us; so that if we shall deal falsely with our God in this work . . . [and] cause him to withdraw his present help from us, we shall be made a story and a byword through the world. byword:example According to Winthrop, why was it important for the Puritans to behave as they believed God would want them to? Choices: [(a) because France might attack the colony if it was weak, (b) because the colonists were not allowed to go back to England, (c) because the whole world was watching, (d) because the colonists could become wealthy] Answer index: 2
@@
Subject: Social Studies Skill: the-constitutional-convention Description: Before the states adopted the Constitution, the powers of the central government were described in the Articles of Confederation. The Articles of Confederation were written quickly and accepted by the states in the middle of the Revolutionary War. The committee that wrote the Articles created a central government that could organize the war. However, the committee did not trust an overly powerful government, so they did not give the government the power to enforce, or make people follow, the law. Overall, they gave the central government very few powers. Under the Articles of Confederation, which of the following powers did the central government have? Choices: [(a) forcing states to contribute money through taxes, (b) forcing states to provide troops, (c) enforcing laws, (d) negotiating treaties with other countries] Answer index: 3
Subject: Social Studies Skill: the-civil-war-the-conclusion-of-the-war Description: In late 1864, a Union army under the command of General William Tecumseh Sherman marched through the state of Georgia. The passage below comes from a letter that Sherman wrote to the War Department about his march. Read the passage. Then answer the question below. We are not only fighting hostile armies, but a hostile people, and must make old and young, rich and poor, feel the hard hand of war, as well as their organized armies. I know that this recent movement of mine through Georgia has had a wonderful effect in this respect. people:the civilians, or people not in the military, of the Confederacy What was Sherman arguing in this passage? Choices: [(a) Most white people in the Confederacy had always supported the Union., (b) The war should only be fought against Confederate soldiers, not civilians., (c) The Union army should make life hard for Confederate civilians., (d) Sherman’s march through Georgia had been a failure for the Union army.] Answer index: 2
Subject: Social Studies Skill: trade-and-specialization Description: Adam Smith was an important thinker from the 1700s whose ideas about trade influenced many people. The passage below comes from one of his famous books. Read the passage. Then answer the question below. It is a [rule] of every [wise] master of a family, never to attempt to make at home what it will cost him more to make than to buy. The tailor does not attempt to make his own shoes, but buys them [from] the shoemaker. The shoemaker does not attempt to make his own clothes but employs a tailor. Adam Smith, The Wealth of Nations According to Adam Smith, how do people benefit from specialization and trade? Choices: [(a) People make clothes and shoes at home and listen to the master of the family., (b) People spend all of their money to buy things they need., (c) People buy things they want or need for less than it would cost to make things themselves., (d) People make everything they need at home to save as much money as possible.] Answer index: 2
@@
Subject: Social Studies Skill: the-civil-war-gettysburg-to-the-election-of-1864 Description: The goal of the Union blockade was to stop the Confederacy from trading with other countries. Even before the war, Northerners and Southerners predicted what would happen if the South’s trade was blockaded. The passage below comes from an 1858 speech by the Southern senator James Henry Hammond. Read the passage. Then answer the question below. Without firing a gun, without drawing a sword, should [the North] make war on us, we could bring the whole world to our feet. . . . What would happen if no cotton was furnished for three years? . . . England would topple headlong and carry the whole civilized world with her, to save the South. No, you dare not to make war on cotton. No power on the earth dares to make war upon it. Cotton is king. England:the United Kingdom According to Hammond, what made the South powerful? Choices: [(a) The South controlled the best harbors in North America., (b) The South had a large enslaved population., (c) The South controlled the world’s cotton supply., (d) The South had given the United Kingdom a large loan.] Answer index: 2
Subject: Social Studies Skill: the-crusades Description: Historians do not have a copy of Pope Urban’s 1095 speech. However, some of the people who heard his speech wrote down what they remembered. Read part of the speech that one person remembered hearing. Then answer the question below. All who die [on their way to Jerusalem], whether by land or by sea, or in battle . . . shall have immediate remission of sins. Jacques Bongars, Gesta Dei per Francos, trans in Oliver J. Thatcher and Edgar Holmes McNeal, eds., A Source Book for Medieval History. Copyright 1905 by Scribners. According to the speech, what did Pope Urban say would happen to people who died while traveling to Jerusalem on a crusade? Choices: [(a) Their land would be taken by the government., (b) All of their bad actions would be forgiven., (c) They would be remembered forever., (d) Their families would be given money.] Answer index: 1
@@
Subject: Social Studies Skill: medieval-japan Description: Between the years 700 and 1100 CE, several families became increasingly wealthy and powerful. Read the passage about how these families became powerful. Then answer the question below. In the capital city, Kyoto, some wealthy families would arrange for their daughters to marry sons from the royal family. Marriages allowed these wealthy families to influence some of the emperor’s decisions. These same families often controlled large amounts of land. Around the year 1100, almost half the land in Japan was controlled by wealthy families! In order to protect their land and wealth, these families often hired their own armies. Eventually, these armies became larger and more powerful than the emperor’s own troops. Based on the passage, how had the status of powerful families changed in Japan by the year 1100? Choices: [(a) By the year 1100, the emperor was no longer the most powerful person in Japan. As a result, he was not able to control the powerful families., (b) Powerful families regularly married their daughters to members of the royal family. By the year 1100, these marriages had helped the emperor become even more powerful., (c) By 1100, most powerful families had become less wealthy. As a result, powerful families were no longer able to pay their private armies.] Answer index: 0
Subject: Social Studies Skill: the-songhai-empire Description: Sunni Ali Ber and his army conquered many territories to build the Songhai Empire. But not everyone in the new empire was happy under his rule. For example, many Muslims, or followers of the religion of Islam, were unhappy. Read the passage by a Muslim scholar who lived in the Songhai Empire during the reign of Sunni Ali. Then answer the question below. This villain,—I mean, [Sunni Ali],—was a prince, who was obeyed and respected. He was feared by his subjects because of his cruelty. . . . The ferocity of [Sunni] Ali towards Muslims and his cruelty . . . in matters of religion, made the hearts of the people overflow with sadness and worry. Christopher Wise, translator, Ta’rikh al fattash. Copyright 2011 by Africa World Press. According to the author, why were Muslims unhappy with Sunni Ali? Choices: [(a) He was a prince., (b) He was a devoted Muslim., (c) He was cruel to Muslims., (d) He was greedy.] Answer index: 2

### v3

@@ Appendix D Summary of Skills
Subject: Social Studies
Skill: causes-of-the-american-revolution-introduction-to-mercantilism-and-the-navigation-acts
Description: To increase Great Britain’s wealth, Parliament passed laws that gave each part of the British Empire a different economic role:
Colonists would send raw materials to people in Great Britain. Raw materials, such as raw iron and wool, are directly from nature.
People in Great Britain would turn the raw materials into manufactured goods, or finished products, such as fabric and tools. British merchants could sell these manufactured goods back to the colonies or to other countries.
Complete the text.
Parlaiment wanted the country to earn more money from selling exports than it spent buying imports. The raw materials that Great Britain imported from the colonies were ___ the manufactured goods that Great Britain exported for colonists to purchase. Parliament’s policies were part of what created the system of Triangular Trade.
Choices: [(a) more valuable than, (b) as valuable as, (c) less valuable than]
Answer index: 2
Subject: Social Studies
Skill: the-revolutionary-war-preparing-for-war
Description: After 1763, the British Parliament began to tax the Thirteen Colonies more heavily to help pay off debt from the French and Indian War. Parliament also restricted colonial trade. Many colonists resented these new taxes and restrictions.
After the passage of the Stamp Act in 1765, which required Americans to pay a tax on all paper they used, an American colonist drew a cartoon with this caption:
The moral is, that the colonies may be ruined, but that Britain would thereby be maimed.
maimed:permanently wounded
What is the author most likely arguing in the caption?
Choices: [(a) The colonies are not a part of the British Empire., (b) Great Britain is harming itself by taxing the colonies., (c) Great Britain’s taxation policies are good for the colonies., (d) The colonies will be better off if they declare independence from Great Britain.]
Answer index: 1
Subject: Social Studies
Skill: u-s-mexican-war
Description: By the 1840s, belief in Manifest Destiny had become quite popular in the United States. So, support for expanding the territory of the U.S. was an important issue in the 1844 presidential election. This election was won by James K. Polk, the candidate of the Democratic Party.
Read the following passage adapted from the 1844 Democratic Party platform, or list of the party’s goals. Then complete the sentence below.
Our title to the whole of the Territory of Oregon is clear and unquestionable, and the annexation of Texas to the United States at the earliest moment possible is important to American progress.
annexation:addition
Complete the sentence.
According to the platform, the Democratic Party ___ Manifest Destiny.
Choices: [(a) supported, (b) opposed, (c) neither supported nor opposed]
Answer index: 0
Subject: Social Studies
Skill: the-constitution-amendments
Description: When the Constitution was written, many Americans were afraid that it made the federal government too powerful. So, ten amendments were added to the Constitution to describe and protect the rights of individual people. These first ten amendments are called the Bill of Rights.
The following passage is from a letter written by Thomas Jefferson to James Madison in 1787, before the Bill of Rights had been approved. Read the passage. Then answer the question below.
A bill of rights is what the people are entitled to . . . and [it is] what no just government should refuse [to write down].
just:fair
Based on the passage, what was Jefferson’s attitude toward a bill of rights?
Choices: [(a) Governments should protect people’s rights, but those rights don’t need to be written down., (b) Writing down a bill of rights doesn’t do anything, because governments can just ignore it., (c) People deserve to have a bill of rights that describes and protects their individual rights.]
Answer index: 2
Subject: Social Studies
Skill: the-revolutionary-war-struggle-for-independence
@@
Description: As the Continental Army fought the British, the Second Continental Congress sent an Olive Branch Petition to King George III in July 1775. The petition stated that the colonies would stop fighting with Great Britain under certain conditions. Which of the following was one of those conditions?
Choices: [(a) if a new colony was established on the western frontier, (b) if Great Britain repealed certain laws and taxes, (c) if the colonies could have their own king and Parliament, (d) if King George III stepped down and a new king was named]
Answer index: 1
Subject: Social Studies
Skill: the-end-of-reconstruction-and-the-rise-of-jim-crow
Description: Some white Southerners believed deeply that they should be shown more respect than African American people. They did not like the Republicans’ plans to protect African American people’s rights. Sometimes their anger at the changes to Southern society turned violent.
The passage below comes from an investigation into the murder of an African American man named Johnson Stuart. The man’s brother is describing what happened before the murder. Read the passage. Then answer the question below.
Question—Did you ever hear of any threats being made against him before he was killed? . . .
State of South Carolina, Evidence taken by the Committee of Investigation of the Third Congressional District, 1870.
According to this passage, why did some white Southerners want to kill Johnson Stuart?
Choices: [(a) They thought he disrespected white people by treating all the voters the same., (b) They thought he was cheating by paying people to vote for him for sheriff., (c) They thought he had stolen money and supplies from the sheriff’s office in the courthouse., (d) They thought he was violating the Constitution by not allowing African Americans to vote.]
Answer index: 0
Subject: Social Studies
Skill: pre-contact-native-peoples-introduction
Description: For many years, most archaeologists agreed that the first humans who came to North America walked across the land bridge around 12,000 years ago. However, new evidence has been discovered that makes some archaeologists believe people may have traveled in a different way.
The text below describes two pieces of evidence that archaeologists use to support a different theory of migration to North America. Look at the text. Then complete the sentence below.
Hunter-gatherers lived in the islands of Japan 20,000 years ago. They used resources from the ocean to survive.
People on the coast of North America used small boats to travel along the coast more than 10,000 years ago.
The text below describes a new theory about how humans arrived in North America. Use the evidence above to complete the text.
According to the new theory, the first humans may have traveled from Asia to North America ___. They may have ___ along the coast of the land bridge.
Choices: [(a) 100,000 years ago . . . walked, (b) using small boats . . . sailed]
Answer index: 1
Subject: Social Studies
Skill: the-reformation
Description: Read two of Martin Luther’s statements criticizing the Catholic Church. Then answer the question below.
It is certain that when money clinks in the [Catholic Church’s] money chest, greed…can be increased.
William R. Russell, translator, The Ninety-Five Theses and Other Writings. Copyright 2017 by Penguin Classics.
Based on the statements, what was one reason why many people were unhappy with the Catholic Church?
Choices: [(a) They thought that the Catholic Church had become too big., (b) They felt their religious community was too small., (c) They believed the Catholic Church should collect more money., (d) They thought the Catholic Church was too rich and greedy.]
Answer index: 3
Subject: Social Studies
Skill: rights-and-responsibilities-of-active-citizenship
Description: American presidents often speak about the rights, duties, and responsibilities of American citizens. Read the quotations from two former presidents. Then answer the question below.
Ask not what your country can do for you; ask what you can do for your country.
Ronald Reagan, Veterans Day Proclamation, October 26, 1981
Based on the quotations above, which of the following statements would both presidents agree with?
Choices: [(a) The country should not expect citizens to do things for their communities., (b) Citizens should expect the government to do everything for them., (c) Active and engaged citizens are an important part of American democracy.]
Answer index: 2
Subject: Social Studies
Skill: the-age-of-exploration-origins
@@
Description: In the 1400s, Europe was made up of many independent states. These states often competed with each other for resources, including imported goods from Africa and Asia. Imagine you are a European ruler in the 1400s. What could you do as a ruler to gain a competitive advantage, or an upper hand, over other European states?
Choices: [(a) try to find new trade routes to Africa and Asia, (b) ban any goods that came from Africa or Asia, (c) rely on other European states to sell goods from Africa and Asia to you]
Answer index: 0
@@
Subject: Social Studies
Skill: industrialization-in-the-gilded-age
Description: Industrialization brought dramatic changes to Americans’ daily lives. Read a modern scholar’s description of those changes. Then answer the question below.
@@
Within a few decades, urban American homes became networked. . . . Instead of relying on candles and [fuel] carried into the home, each home was connected to the electricity network that provided electric light . . . Instead of relying on . . . outhouses and cesspools, each home was gradually connected to two more networks, one bringing in a supply of clean running water and the other taking waste out into sewers. Houses of the rich after 1880 and of the working class after 1910 were increasingly supplied with central heating.
Robert J. Gordon, The Rise and Fall of American Growth, Princeton University Press, 2016.
What is the main idea of the passage?
Choices: [(a) As the U.S. industrialized, fewer homes had electricity and central heating., (b) Industrialization gave many Americans much more comfortable lives., (c) Electric lighting and sewers made homes dirtier and more dangerous., (d) Factories became very modern, but regular people’s houses did not change.]
Answer index: 1
Subject: Social Studies
Skill: causes-of-the-american-revolution-politics-and-society
@@
Description: Andrew Burnaby, an English minister, traveled around the Thirteen Colonies during the French and Indian War. Based on his observations, Burnaby argued that America would never be independent of Great Britain. Read Burnaby’s argument. Then answer the question below.
The difference of character, of manners, of religion, of interest, of the different colonies [is so great that if they were] left to themselves there would soon be a civil war.
Why did Burnaby think that America would never be a nation independent from Great Britain?
Choices: [(a) The colonies were united in their support of Great Britain., (b) Colonial businesses needed British customers., (c) There were too many differences between the colonies., (d) The colonies signed a pledge to Great Britain after the French and Indian War.]
Answer index: 2
Subject: Social Studies
Skill: new-england-colonies-founding-and-government
Description: Members of a religious group called the Puritans founded the Massachusetts Bay Colony. Unlike the Pilgrims, the Puritans did not want to leave the Church of England. Instead, they wanted to change the Church of England according to their religious beliefs.
The first governor of the Massachusetts Bay Colony was a Puritan named John Winthrop. The passage below comes from a speech that Winthrop gave to the Puritans on their voyage to New England. Read the passage. Then answer the question below.
For we must consider that we shall be as a city upon a hill. The eyes of all people are upon us; so that if we shall deal falsely with our God in this work . . . [and] cause him to withdraw his present help from us, we shall be made a story and a byword through the world.
byword:example
According to Winthrop, why was it important for the Puritans to behave as they believed God would want them to?
Choices: [(a) because France might attack the colony if it was weak, (b) because the colonists were not allowed to go back to England, (c) because the whole world was watching, (d) because the colonists could become wealthy]
Answer index: 2
Subject: Social Studies
Skill: the-constitutional-convention
Description: Before the states adopted the Constitution, the powers of the central government were described in the Articles of Confederation.
@@
The Articles of Confederation were written quickly and accepted by the states in the middle of the Revolutionary War. The committee that wrote the Articles created a central government that could organize the war. However, the committee did not trust an overly powerful government, so they did not give the government the power to enforce, or make people follow, the law. Overall, they gave the central government very few powers.
Under the Articles of Confederation, which of the following powers did the central government have?
Choices: [(a) forcing states to contribute money through taxes, (b) forcing states to provide troops, (c) enforcing laws, (d) negotiating treaties with other countries]
Answer index: 3
Subject: Social Studies
Skill: the-civil-war-the-conclusion-of-the-war
Description: In late 1864, a Union army under the command of General William Tecumseh Sherman marched through the state of Georgia. The passage below comes from a letter that Sherman wrote to the War Department about his march. Read the passage. Then answer the question below.
We are not only fighting hostile armies, but a hostile people, and must make old and young, rich and poor, feel the hard hand of war, as well as their organized armies. I know that this recent movement of mine through Georgia has had a wonderful effect in this respect.
people:the civilians, or people not in the military, of the Confederacy
What was Sherman arguing in this passage?
Choices: [(a) Most white people in the Confederacy had always supported the Union., (b) The war should only be fought against Confederate soldiers, not civilians., (c) The Union army should make life hard for Confederate civilians., (d) Sherman’s march through Georgia had been a failure for the Union army.]
Answer index: 2
Subject: Social Studies
Skill: trade-and-specialization
Description: Adam Smith was an important thinker from the 1700s whose ideas about trade influenced many people. The passage below comes from one of his famous books. Read the passage. Then answer the question below.
It is a [rule] of every [wise] master of a family, never to attempt to make at home what it will cost him more to make than to buy. The tailor does not attempt to make his own shoes, but buys them [from] the shoemaker. The shoemaker does not attempt to make his own clothes but employs a tailor.
Adam Smith, The Wealth of Nations
According to Adam Smith, how do people benefit from specialization and trade?
Choices: [(a) People make clothes and shoes at home and listen to the master of the family., (b) People spend all of their money to buy things they need., (c) People buy things they want or need for less than it would cost to make things themselves., (d) People make everything they need at home to save as much money as possible.]
Answer index: 2
Subject: Social Studies
Skill: the-civil-war-gettysburg-to-the-election-of-1864
@@
Description: The goal of the Union blockade was to stop the Confederacy from trading with other countries. Even before the war, Northerners and Southerners predicted what would happen if the South’s trade was blockaded. The passage below comes from an 1858 speech by the Southern senator James Henry Hammond. Read the passage. Then answer the question below.
Without firing a gun, without drawing a sword, should [the North] make war on us, we could bring the whole world to our feet. . . . What would happen if no cotton was furnished for three years? . . . England would topple headlong and carry the whole civilized world with her, to save the South. No, you dare not to make war on cotton. No power on the earth dares to make war upon it. Cotton is king.
England:the United Kingdom
According to Hammond, what made the South powerful?
Choices: [(a) The South controlled the best harbors in North America., (b) The South had a large enslaved population., (c) The South controlled the world’s cotton supply., (d) The South had given the United Kingdom a large loan.]
Answer index: 2
Subject: Social Studies
Skill: the-crusades
Description: Historians do not have a copy of Pope Urban’s 1095 speech. However, some of the people who heard his speech wrote down what they remembered. Read part of the speech that one person remembered hearing. Then answer the question below.
All who die [on their way to Jerusalem], whether by land or by sea, or in battle . . . shall have immediate remission of sins.
Jacques Bongars, Gesta Dei per Francos, trans in Oliver J. Thatcher and Edgar Holmes McNeal, eds., A Source Book for Medieval History. Copyright 1905 by Scribners.
According to the speech, what did Pope Urban say would happen to people who died while traveling to Jerusalem on a crusade?
Choices: [(a) Their land would be taken by the government., (b) All of their bad actions would be forgiven., (c) They would be remembered forever., (d) Their families would be given money.]
Answer index: 1
Subject: Social Studies
Skill: medieval-japan
Description: Between the years 700 and 1100 CE, several families became increasingly wealthy and powerful. Read the passage about how these families became powerful. Then answer the question below.
@@
In the capital city, Kyoto, some wealthy families would arrange for their daughters to marry sons from the royal family. Marriages allowed these wealthy families to influence some of the emperor’s decisions. These same families often controlled large amounts of land. Around the year 1100, almost half the land in Japan was controlled by wealthy families! In order to protect their land and wealth, these families often hired their own armies. Eventually, these armies became larger and more powerful than the emperor’s own troops.
Based on the passage, how had the status of powerful families changed in Japan by the year 1100?
Choices: [(a) By the year 1100, the emperor was no longer the most powerful person in Japan. As a result, he was not able to control the powerful families., (b) Powerful families regularly married their daughters to members of the royal family. By the year 1100, these marriages had helped the emperor become even more powerful., (c) By 1100, most powerful families had become less wealthy. As a result, powerful families were no longer able to pay their private armies.]
Answer index: 0
Subject: Social Studies
Skill: the-songhai-empire
Description: Sunni Ali Ber and his army conquered many territories to build the Songhai Empire. But not everyone in the new empire was happy under his rule. For example, many Muslims, or followers of the religion of Islam, were unhappy.
Read the passage by a Muslim scholar who lived in the Songhai Empire during the reign of Sunni Ali. Then answer the question below.
This villain,—I mean, [Sunni Ali],—was a prince, who was obeyed and respected. He was feared by his subjects because of his cruelty. . . . The ferocity of [Sunni] Ali towards Muslims and his cruelty . . . in matters of religion, made the hearts of the people overflow with sadness and worry.
Christopher Wise, translator, Ta’rikh al fattash. Copyright 2011 by Africa World Press.
According to the author, why were Muslims unhappy with Sunni Ali?
Choices: [(a) He was a prince., (b) He was a devoted Muslim., (c) He was cruel to Muslims., (d) He was greedy.]
Answer index: 2
@@
Subject: Social Studies
Skill: origins-of-hinduism